Let a be an element of.order n in a group and k be a positive integer .prove that |ak| = n/(gcd (n,k))


1
Expert's answer
2021-08-05T17:15:03-0400

Let "a" be an element of order "n" in a group "G" with identity "e," and "k" be a positive integer. Let "n=dn_1, k=dk_1," where "d=gcd(n,k)." Then "(n_1, k_1)=1." Let "|a^k|=t." If "(a^k)^t=e," then "a^{kt}=e," and hence "n=dn_1" is a divisor of "kt=dk_1t." Since "(n_1, k_1)=1," we conclude that "n_1" divides "t." Taking into account that "(a^k)^{n_1}=a^{dk_1n_1}=a^{nk_1}=(a^n)^{k_1}=e" and "n_1" divides "t," we conclude that "t=n_1," and hence "|a^k|=n_1=\frac{n}{d}=\frac{n}{gcd(n,k)}."